Find volume bounded by the planes 𝑥 = 0, 𝑦 = 0, 𝑧 = 0 and 2𝑥 + 𝑦 + 𝑧 = 6 0 ≤ 𝑥 ≤ 2,0 ≤ 𝑦 ≤ 6 − 2�



1
Expert's answer
2021-06-11T02:59:30-0400
"V=\displaystyle\int_{0}^{2}\displaystyle\int_{0}^{6-2x}\displaystyle\int_{0}^{6-y-2x}dzdydx"

"=\displaystyle\int_{0}^{2}\displaystyle\int_{0}^{6-2x}[z]\begin{matrix}
 6-y-2x \\
 0
\end{matrix}dydx"

"=\displaystyle\int_{0}^{2}\displaystyle\int_{0}^{6-2x}(6-y-2x)dydx"

"=\displaystyle\int_{0}^{2}[6y-\dfrac{y^2}{2}-2xy]\begin{matrix}
 6-2x \\
 0
\end{matrix}dx"

"=\displaystyle\int_{0}^{2}(36-12x-18+12x-2x^2-12x+4x^2)dx"

"=\displaystyle\int_{0}^{2}(2x^2-12x+18)dx"


"=[\dfrac{2}{3}x^3-6x^2+18x]\begin{matrix}
 2 \\
 0
\end{matrix}"

"=\dfrac{16}{3}-24+36=\dfrac{52}{3}(units^3)"

"V=\dfrac{52}{3}" units of volume.